### v2

# Efficient Subdivision of Bézier Curves/Surfaces via Blossoms

Abstract
We consider the problem of Bézier curves/surfaces subdivision using blossoms. We propose closed-form formulae for blossoms evaluation, as needed for the calculation of control points. This approach leads to direct and efficient way to obtain subdivisions for Bézier curves and both tensor product and triangular Bézier surfaces. It simplifies considerably the computation of control points of subdivisions which is crucial in applications where curves/surfaces need to be refined or adapted dynamically. For instance, in CAD/CAM systems, architectural design, or animation, the ability to quickly and accurately determine new control points is essential for manipulation and rendering complex shapes. More efficient subdivision can facilitate complex operations like finding intersections between surfaces or smoothly blending multiple surfaces.
keywords:
Bézier curve , tensor product Bézier surface , triangular Bézier surface , blossom , subdivision
MSC:
65D17 , 65D18 , 65D19 , 68U05 , 68U07
journal: arXiv.org
[…]
[inst1]organization=Faculty of Mathematics and Informatics, Sofia University “St. Kliment Ohridski”, addressline=Blvd. James Bourchier 5, city=Sofia, postcode=1164, country=Bulgaria
1 Introduction
Blossoming is a fundamental tool for analyzing polynomial and piecewise polynomial curves and surfaces. It enables derivation of important identities and algorithms (Goldman, 2003; Mann, 2006). It was originally introduced to the field of Computer Aided Geometric Design (CAGD) by de Casteljau (de Casteljau, 1985) and later developed by Ramshaw (Ramshaw, 1987, 1989) for the study of Bézier and B-splines curves and surfaces.
The blossom of a polynomial $P(t)$ of degree $n$ is the unique multivariate polynomial $p=p(u_{1},\dots,u_{n})$ satisfying the following properties (Goldman, 2003):
  1. (i)
    Symmetry. Polynomial $p$ is invariant under permutation of its parameters, i.e. $p(u_{\sigma(1)},\dots,u_{\sigma(n)})=p(u_{1},\dots,u_{n})$ for any permutation $\sigma$ of $\{1,\dots,n\}$.
  2. (ii)
    Multi-affinity. Polynomial $p$ is affine in each of its variables i.e.
    $p(\dots,(1-\alpha)u_{k}+\alpha v_{k},\dots)=(1-\alpha)p(\dots,u_{k},\dots)+\alpha p(\dots,v_{k},\dots).$
  3. (iii)
    Diagonal reduction. Polynomial $p$ reduces to the original polynomial $P(t)$ along the diagonal, i.e. $p(t,\dots,t)=P(t).$
The definition of blossom in the bivariate case is extended straightforward (Goldman, 2003).
Blossoming approach offers significant utility in various aspects of CAGD theory including curve and surface manipulation, computing derivatives and identities, establishing smoothness conditions, changing bases, dual functionals and others, see (Goldman, 2003; Mann, 2006). In CAD systems blossoming facilitates evaluation and manipulation of control polygons and nets for subdivisions, enables high precision geometric representations of complex surfaces, supports smooth animation via rapid computation and interpolation of control points, enhances control over surface continuity and adaptability, enabling the design of intricate shapes in various engineering fields and others.
[…]
In this paper we consider the following three problems:
  1. (1)
    Given a polynomial curve $C(u)=\sum_{i=0}^{n}{\bf c}_{i}u^{i}$ of degree $n$, $n\in\mathbb{N}$, ${\bf c}_{i}\in\mathbb{R}^{3}$, $i=0,\dots,n$, find Bézier curve $C^{1}$ on $C$ defined for $a\leq u\leq b$, where $a,b\in\mathbb{R}$;
  2. (2)
    Given a polynomial surface $S(u,v)=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}u^{i}v^{j}$ of degree $(n,m)$, $n,m\in\mathbb{N}$, ${\bf c}_{ij}\in\mathbb{R}^{3}$, $i=0,\dots,n$, $j=0,\dots,m$, find tensor product Bézier (TPB) surface patch $S^{1}$ on $S$ defined for $a\leq u\leq b$ and $c\leq v\leq d$, where $a,b,c,d\in\mathbb{R}$;
  3. (3)
    Given a polynomial surface $S(u,v)=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}u^{i}v^{j}$ of degree $(n,m)$, $n,m\in\mathbb{R}^{3}$, ${\bf c}_{ij}\in\mathbb{N}$, $i=0,\dots,n$, $j=0,\dots,m$, find triangular Bézier (TB) surface patch $S^{2}$ on $S$ defined in a triangle $\triangle\bf{abc}$, where ${\bf a},{\bf b},{\bf c}\in\mathbb{R}^{2}$.
Standard approach to derive the Bézier curve $C^{1}$ and the Bézier surface patches $S^{1}$ and $S^{2}$ is subdivision using blossoming as it is detailed in Section 2. The control points of $C^{1}$, $S^{1}$, and $S^{2}$ are expressed in terms of the values of the corresponding blossoms. They solely depend on the parameters $a,b$; $a,b,c,d$; $\bf{a},\bf{b},\bf{c}$, respectively. However, the relevant expressions for the control points are long and computationally intractable sums. We propose a closed-form formulae for computation of the control points for each of the above three problems. These formulae provide a straightforward and effective procedure for computation of control points and facilitate the evaluation process for the three considered problems.
The paper is organized as follows: In Section 2 we describe the solutions of the considered problems by blossoming. Then we present closed-form formulae for computation of the control points and prove their correctness. In Section 3 we present and discuss results from the numerical implementation and visualization of the closed-form formulae. At the end we provide some concluding remarks.
2 Main Results
2.1 Subdivision of a Bézier Curve
Let $C(u)=\sum_{i=0}^{n}{\bf c}_{i}u^{i}$ be a polynomial curve of degree $n$, $n\in\mathbb{N}$, ${\bf c}_{i}\in\mathbb{R}^{3}$, $i=0,\dots,n$. Let the Bézier curve $C^{1}(u)=\sum_{\nu=0}^{n}{\bf w}_{\nu}B_{\nu}^{n}(u)$ be the part of $C(u)$ defined in the interval $[a,b]$, where $a,b\in\mathbb{R}$, ${\bf w}_{\nu}$ are the control points of $C^{1}$, and $B_{\nu}^{n}(u)=\binom{n}{\nu}u^{\nu}(1-u)^{n-\nu}$ are the Bernstein polynomials, $\nu=0,\dots,n$.
Let $b\tilde{\ }(u_{1},\dots,u_{n})$ be the blossom of $C(u)$. It is well known that
${\bf w}_{\nu}=b\tilde{\ }(\underbrace{b,\dots,b}_{\text{$\nu$}},\underbrace{a,\dots,a}_{\text{$n-\nu$}}).$ (1)
Furthermore, for the blossom $b\tilde{\ }$ we have the formula
$b\tilde{\ }(u_{1},\dots,u_{n})=\sum_{i=0}^{n}{\bf c}_{i}b_{i}\tilde{\ }(u_{1},\dots,u_{n}),\ {\rm where}$
$b_{i}\tilde{\ }(u_{1},\dots,u_{n})=\sum_{\{\alpha_{1},\dots,\alpha_{i}\}\subseteq\{1,\dots,n\}}\frac{u_{\alpha_{1}}\dots u_{\alpha_{i}}}{\binom{n}{i}}$ (2)
is the blossom of the monomial $u^{i}$. The sums in (2) are taken over all distinct indices $\{\alpha_{1},\dots\alpha_{i}\}$ and $b_{0}\tilde{\ }(u_{1},\dots,u_{n})=1$.
We have observed that the evaluation of blossoms using formula (2) is not straightforward since it requires the generation of all subsets $\{\alpha_{1},\dots,\alpha_{i}\}$. This make the evaluation of control points computationally intense. In the next theorem we propose a formula for direct computation of the control points ${\bf w}_{\nu}$, $\nu=0,\dots,n$.
Theorem 1
The Bézier control points ${\bf w}_{\nu}$, $\nu=0,\dots,n$, defined by (1) are
${\bf w}_{\nu}=\sum_{i=0}^{n}\frac{{\bf c}_{i}}{\binom{n}{i}}\sum_{k=\max(0,i+\nu-n)}^{\min(i,\nu)}\binom{\nu}{k}\binom{n-\nu}{i-k}b^{k}a^{i-k}.$ (3)
Proof According to (2), we have
$b_{i}\tilde{\ }(u_{1},\dots,u_{n}):=\frac{1}{\binom{n}{i}}\sum_{\bar{\alpha}_{i}\in{\cal A}_{i}}U(\bar{\alpha}_{i})$
where ${\cal A}_{i}$ is the set of all subsets $\bar{\alpha}_{i}:=\{\alpha_{1},\dots,\alpha_{i}\}$ of distinct indices of size $i$ and $U(\bar{\alpha}_{i}):=u_{\alpha_{1}}\dots u_{\alpha_{i}}$.
Next, we will compute $b_{i}\tilde{\ }(\underbrace{b,\dots,b}_{\text{$\nu$}},\underbrace{a,\dots,a}_{\text{$n-\nu$}})$. For a fixed $\bar{\alpha}_{i}$ let $k=k(\bar{\alpha}_{i})$ be the number of indices in $\bar{\alpha}_{i}$ that are in $1,\nu]$. Then the number of indices in $\bar{\alpha}_{i}$ that are in $\nu+1,n]$ is iki-k. Note that the following inequalities must hold
$k\leq i,\ k\leq\nu,\ i-k\leq n-\nu.$ (4)
In this notation for a fixed $\bar{\alpha}_{i}$ we have $U(\bar{\alpha}_{i})=b^{k}a^{i-k}$. Consequently,
$\sum_{\bar{\alpha}_{i}\in{\cal A}_{i}}U(\bar{\alpha}_{i})=\sum_{k=0}^{i}I(k)b^{k}a^{i-k}$
where $I(k)$ is the number of subsets $\bar{\alpha}_{i}$ that have $k$ indices in $1,\nu]$. We have $I(k)=\binom{\nu}{k}\binom{n-\nu}{i-k}$ and hence, from (4) it follows
$\sum_{\bar{\alpha}_{i}\in{\cal A}_{i}}U(\bar{\alpha}_{i})=\sum_{k=\max(0,i+\nu-n)}^{\min(i,\nu)}\binom{\nu}{k}\binom{n-\nu}{i-k}b^{k}a^{n-k}.$ (5)
Substituting (5) in (2) and (1), we get
${\bf w}_{\nu}=\sum_{i=0}^{n}\frac{{\bf c}_{i}}{\binom{n}{i}}\sum_{k=\max(0,i+\nu-n)}^{\min(i,\nu)}\binom{\nu}{k}\binom{n-\nu}{i-k}b^{k}a^{i-k}.$
$\Box$
2.2 Subdivision of a Tensor Product Bézier Surface
Let $S(u,v)=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}u^{i}v^{j}$ be a polynomial surface of bidegree $(n,m)$, $n,m\in\mathbb{N}$, ${\bf c}_{ij}\in\mathbb{R}^{3}$, $i=0,\dots,n$, $j=0,\dots,m$. Let the TPB surface $S^{1}(u,v)=\sum_{\nu=0}^{n}\sum_{\mu=0}^{m}{\bf p}_{\nu\mu}B_{\nu}^{n}(u)B_{\mu}^{n}(v)$ be the part of $S(u,v)$ defined for $a\leq u\leq b$ and $c\leq v\leq d$, where $a,b,c,d\in\mathbb{R}$ and ${\bf p}_{\nu\mu}$ are the control points of $S^{1}$, $\nu=0,\dots,n$, $\mu=0,\dots,m$.
Let $b^{\Box}(u_{1},\dots,u_{n},v_{1},\dots,v_{m})$ be the blossom of $S(u,v)$. It is known that
${\bf p}_{\nu\mu}=b^{\Box}(\underbrace{b,\dots,b}_{\text{$\nu$}},\underbrace{a,\dots,a}_{\text{$n-\nu$}},\underbrace{d,\dots,d}_{\text{$\mu$}},\underbrace{c,\dots,c}_{\text{$m-\mu$}}).$ (6)
Furthermore, for the blossom $b^{\Box}$ we have the formula (see (Goldman, 2003), pp. 338-339)
$b^{\Box}(u_{1},\dots,u_{n},v_{1},\dots,v_{m})=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}b_{ij}^{\Box}(u_{1},\dots,u_{n},v_{1},\dots,v_{m}),\ \rm{where}$ (7)
$b_{ij}^{\Box}(u_{1},\dots,u_{n},v_{1},\dots,v_{m})=\sum_{\{\alpha_{1},\dots,\alpha_{i}\}\subseteq\{1,\dots,n\}}\frac{u_{\alpha_{1}}\dots u_{\alpha_{i}}}{\binom{n}{i}}\sum_{\{\beta_{1},\dots,\beta_{j}\}\subseteq\{1,\dots,m\}}\frac{v_{\beta_{1}}\dots v_{\beta_{j}}}{\binom{m}{j}}$ (8)
is the blossom of the monomial $u^{i}v^{j}$. The sums in (8) are taken over all distinct indices $\{\alpha_{1},\dots\alpha_{i}\}$ and all distinct indices $\{\beta_{1},\dots\beta_{i}\}$. In the case where either $i=0$, or $j=0$, the corresponding sum in (8) equals 1.
We note that formula (8) in this form is not easy to be implemented. Therefore, next we establish formula (6) in an equivalent closed form that is more suitable for computations.
Theorem 2
The Bézier control points ${\bf p}_{\nu\mu}$, $\nu=0,\dots,n$, $\mu=0,\dots,m$, defined by (6) are
${\bf p}_{\nu\mu}=\sum_{i=0}^{n}\sum_{j=0}^{m}\frac{{\bf c}_{ij}}{\binom{n}{i}\binom{m}{j}}\sum_{k=\max(0,i+\nu-n)}^{\min(i,\nu)}\ \sum_{r=\max(0,j+\mu-m)}^{\min(j,\mu)}\binom{\nu}{k}\binom{n-\nu}{i-k}\binom{\mu}{r}\binom{m-\mu}{j-r}b^{k}a^{i-k}d^{r}c^{j-r}.$ (9)
Proof We observe that formula (8) can be presented in the equivalent form
$b_{ij}^{\Box}(u_{1},\dots,u_{n},v_{1},\dots,v_{m})=b_{i}\tilde{\ }(u_{1},\dots,u_{n})b_{j}\tilde{\ }(v_{1},\dots,v_{m}),$ (10)
where $b_{i}\tilde{\ }$ and $b_{j}\tilde{\ }$ are the blossoms of the monomials $u^{i}$ and $v^{j}$, respectively. From Theorem 1 we have
$b_{i}\tilde{\ }(u_{1},\dots,u_{n})=\frac{1}{\binom{n}{i}}\sum_{k=\max(0,i+\nu-n)}^{\min(i,\nu)}\binom{\nu}{k}\binom{n-\nu}{i-k}b^{k}a^{n-k},$
$b_{j}\tilde{\ }(v_{1},\dots,v_{m})=\frac{1}{\binom{m}{j}}\sum_{r=\max(0,j+\mu-m)}^{\min(j,\mu)}\binom{\mu}{r}\binom{m-\mu}{j-r}d^{r}c^{j-r}.$
Substituting (11) in (10) and (7), we obtain from (6)
${\bf p}_{\nu\mu}=\sum_{i=0}^{n}\sum_{j=0}^{m}\frac{{\bf c}_{ij}}{\binom{n}{i}\binom{m}{j}}\sum_{k=\max(0,i+\nu-n)}^{\min(i,\nu)}\ \sum_{r=\max(0,j+\mu-m)}^{\min(j,\mu)}\binom{\nu}{k}\binom{n-\nu}{i-k}\binom{\mu}{r}\binom{m-\mu}{j-r}b^{k}a^{i-k}d^{r}c^{j-r}.$
$\Box$
2.3 Subdivision of a Triangular Bézier Surface
Let $S(u,v)=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}u^{i}v^{j}$ be a polynomial surface of total degree $N=n+m$, $n,m\in\mathbb{N}$, ${\bf c}_{ij}\in\mathbb{R}^{3}$, $i=0,\dots,n$, $j=0,\dots,m$, and ${\bf a}=(a_{1},a_{2})$, ${\bf b}=(b_{1},b_{2})$, ${\bf c}=(c_{1},c_{2})$ be three non-collinear points in $\mathbb{R}^{2}$. Let the TB surface
$S^{2}(u,v)=\sum_{\nu=0}^{N}\sum_{\mu=0}^{N-\nu}{\bf q}_{\nu\mu}B_{\nu\mu}^{N}(u,v)u^{\nu}v^{\mu}(1-u-v)^{N-\nu-\mu}$
be the part of $S(u,v)$ defined in the triangle $\triangle\bf{abc}$, where ${\bf q}_{\nu\mu}$ are the control points of $S^{2}$,
$B_{\nu\mu}^{N}(u,v)=\binom{N}{\nu,\mu}u^{\nu}v^{\mu}(1-u-v)^{N-\nu-\mu},\ \nu=0,\dots,N,\ \mu=0,\dots,N-\nu$
are the bivariate Bernstein polynomials, and $\binom{N}{\nu,\mu}=\frac{N!}{\nu!\mu!(N-\nu-\mu)!}$.
Let $b^{\triangle}((u_{1},v_{1}),\dots,(u_{N},v_{N}))$ be the blossom of $S(u,v)$. It is known that
${\bf q}_{\nu\mu}=b^{\triangle}(\underbrace{{\bf a},\dots,{\bf a}}_{\nu},\underbrace{{\bf b},\dots,{\bf b}}_{\mu},\underbrace{{\bf c},\dots\bf c}_{N-\nu-\mu}).$ (12)
Furthermore, for the blossom $b^{\triangle}$ we have the formula (see (Goldman, 2003), pp. 338-339)
$b^{\triangle}((u_{1},v_{1}),\dots,(u_{N},v_{N}))=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}b_{ij}^{\triangle}((u_{1},v_{1}),\dots,(u_{N},v_{N})),\ \rm{where}$
$b_{ij}^{\triangle}((u_{1},v_{1}),\dots,(u_{N},v_{N}))=\sum_{\scriptstyle\{\alpha_{1},\dots,\alpha_{i},\beta_{1},\dots,\beta_{j}\}\subseteq\{1,\dots,N\}}\frac{u_{\alpha_{1}}\dots u_{\alpha_{i}}v_{\beta_{1}}\dots v_{\beta_{j}}}{\binom{N}{i,j}}$, (13)
is the blossom of the monomial $u^{i}v^{j}$. The sum in (13) is taken over all sets of distinct indices $\{\alpha_{1},\dots,\alpha_{i},\beta_{1},\dots,\beta_{j}\}$ and $b_{i0}^{\triangle}((u_{1},v_{1}),\dots,(u_{N},v_{N}))=b_{i}\tilde{\ }(u_{1},\dots,u_{N})$, $b_{0j}^{\triangle}((u_{1},v_{1}),\dots,(u_{N},v_{N}))=b_{j}\tilde{\ }(v_{1},\dots,v_{N}).$
We note that the evaluation of blossoms using formula (13) is not straightforward and easy for implementation. Therefore, similar to the case of TPB surfaces, in the next theorem we propose a formula for direct computation of the control points ${\bf q}_{\nu\mu}$, $\nu=0,\dots,N$, $\mu=0,\dots,N-\nu$.
Theorem 3
The Bézier control points ${\bf q}_{\nu\mu}$, $\nu=0,\dots,N$, $\mu=0,\dots,N-\nu$, defined by (12) are
${\bf q}_{\nu\mu}$ $=$ $\sum_{i=0}^{n}\sum_{j=0}^{m}\frac{{\bf c}_{ij}}{\binom{n+m}{i,j}}\sum_{i_{\alpha}=\underline{i}_{\alpha}}^{\bar{i}_{\alpha}}\sum_{i_{\beta}=\underline{i}_{\beta}}^{\bar{i}_{\beta}}\sum_{j_{\alpha}=\underline{j}_{\alpha}}^{\bar{j}_{\alpha}}\sum_{j_{\beta}=\underline{j}_{\beta}}^{\bar{j}_{\beta}}$ (14)
$\binom{\nu}{i_{\alpha}}\binom{\mu}{i_{\beta}}\binom{\lambda}{i_{\gamma}}\binom{\nu-i_{\alpha}}{j_{\alpha}}\binom{\mu-i_{\beta}}{j_{\beta}}\binom{\lambda-i_{\gamma}}{j_{\gamma}}a_{1}^{i_{\alpha}}b_{1}^{i_{\beta}}c_{1}^{i_{\gamma}}a_{2}^{j_{\alpha}}b_{2}^{j_{\beta}}c_{2}^{j_{\gamma}},$
where
$\begin{array}[]{l}\lambda=n+m-\nu-\mu,\ i_{\gamma}=i-i_{\alpha}-i_{\beta},\ j_{\gamma}=j-j_{\alpha}-j_{\beta},\\ \underline{i}_{\alpha}=\max(0,i-\mu-\lambda),\ \bar{i}_{\alpha}=\min(i,\nu),\\ \underline{i}_{\beta}=\max(0,i-i_{\alpha}-\lambda),\ \bar{i}_{\beta}=\min(i-i_{\alpha},\mu),\\ \underline{j}_{\alpha}=\max(0,j-(\mu-i_{\beta})-(\lambda-i_{\gamma})),\ \bar{j}_{\alpha}=\min(j,\nu-i_{\alpha}),\\ \underline{j}_{\beta}=\max(0,j-j_{\alpha}-(\lambda-i_{\gamma})),\ \bar{j}_{\beta}=\min(j-j_{\alpha},\mu-i_{\beta}).\end{array}$
Proof. First, we will compute $b_{ij}^{\triangle}(\underbrace{{\bf a},\dots,{\bf a}}_{\nu},\underbrace{{\bf b},\dots,{\bf b}}_{\mu},\underbrace{{\bf c},\dots\bf c}_{N-\nu-\mu})$. For fixed $\bar{\alpha}_{i}=\{\alpha_{1},\dots,\alpha_{i}\}$ and $\bar{\beta}_{j}=\{\beta_{1},\dots,\beta_{j}\}$ we have
$u_{\alpha_{1}}\dots u_{\alpha_{i}}v_{\beta_{1}}\dots v_{\beta_{j}}=a_{1}^{i_{\alpha}}b_{1}^{i_{\beta}}c_{1}^{i_{\gamma}}a_{2}^{j_{\alpha}}b_{2}^{j_{\beta}}c_{2}^{j_{\gamma}}\quad{\rm where}$
$\begin{array}[]{l}i_{\alpha}\ {\rm is\ the\ number\ of\ indices\ in}\ \bar{\alpha}_{i}\ {\rm that\ are\ in}\ [1,\nu],\\ i_{\beta}\ {\rm is\ the\ number\ of\ indices\ in}\ \bar{\alpha}_{i}\ {\rm that\ are\ in}\ [\nu+1,\nu+\mu],\\ i_{\gamma}:=i-i_{\alpha}-i_{\beta}\ {\rm is\ the\ number\ of\ indices\ in}\ \bar{\alpha}_{i}\ {\rm that\ are\ in}\ [\nu+\mu+1,N],\\ j_{\alpha}\ {\rm is\ the\ number\ of\ indices\ in}\ \bar{\beta}_{j}\ {\rm that\ are\ in}\ [1,\nu],\\ j_{\beta}\ {\rm is\ the\ number\ of\ indices\ in}\ \bar{\beta}_{j}\ {\rm that\ are\ in}\ [\nu+1,\nu+\mu],\\ j_{\gamma}:=j-j_{\alpha}-j_{\beta}\ {\rm is\ the\ number\ of\ indices\ in}\ \bar{\beta}_{j}\ {\rm that\ are\ in}\ [\nu+\mu+1,N].\end{array}$
For a fixed quadruple $(i_{\alpha},i_{\beta},j_{\alpha},j_{\beta})$ let $I(i_{\alpha},i_{\beta},j_{\alpha},j_{\beta})$ be the set of index vectors $(\bar{\alpha}_{i},\bar{\beta}_{j})$ that have prescribed $i_{\alpha},i_{\beta},j_{\alpha},j_{\beta}$ numbers. Then we have
$b_{ij}^{\triangle}(\underbrace{{\bf a},\dots,{\bf a}}_{\nu},\underbrace{{\bf b},\dots,{\bf b}}_{\mu},\underbrace{{\bf c},\dots\bf c}_{N-\nu-\mu})=\sum_{i_{\alpha},i_{\beta},j_{\alpha},j_{\beta}}|I(i_{\alpha},i_{\beta},j_{\alpha},j_{\beta})|\frac{a_{1}^{i_{\alpha}}b_{1}^{i_{\beta}}c_{1}^{i_{\gamma}}a_{2}^{j_{\alpha}}b_{2}^{j_{\beta}}c_{2}^{j_{\gamma}}}{\binom{N}{i,j}}$
where $|I(i_{\alpha},i_{\beta},j_{\alpha},j_{\beta})|$ denotes the cardinality of the set $I(i_{\alpha},i_{\beta},j_{\alpha},j_{\beta})$. We have
$|I(i_{\alpha},i_{\beta},j_{\alpha},j_{\beta})|=\binom{\nu}{i_{\alpha}}\binom{\mu}{i_{\beta}}\binom{N-\nu-\mu}{i_{\gamma}}\binom{\nu-i_{\alpha}}{j_{\alpha}}\binom{\mu-i_{\beta}}{j_{\beta}}\binom{N-\nu-\mu-i_{\gamma}}{j_{\gamma}}$. (15)
Hence,
$b_{ij}^{\triangle}(\underbrace{{\bf a},\dots,{\bf a}}_{\nu},\underbrace{{\bf b},\dots,{\bf b}}_{\mu},\underbrace{{\bf c},\dots\bf c}_{N-\nu-\mu})=$
$\sum_{j_{\beta}=\underline{j}_{\beta}}^{\bar{j}_{\beta}}\sum_{j_{\alpha}=\underline{j}_{\alpha}}^{\bar{j}_{\alpha}}\sum_{i_{\beta}=\underline{i}_{\beta}}^{\bar{i}_{\beta}}\sum_{i_{\alpha}=\underline{i}_{\alpha}}^{\bar{i}_{\alpha}}\binom{\nu}{i_{\alpha}}\binom{\mu}{i_{\beta}}\binom{\lambda}{i_{\gamma}}\binom{\nu-i_{\alpha}}{j_{\alpha}}\binom{\mu-i_{\beta}}{j_{\beta}}\binom{\lambda-i_{\gamma}}{j_{\gamma}}\frac{a_{1}^{i_{\alpha}}b_{1}^{i_{\beta}}c_{1}^{i_{\gamma}}a_{2}^{j_{\alpha}}b_{2}^{j_{\beta}}c_{2}^{j_{\gamma}}}{\binom{N}{i,j}}$,
where $\lambda:=N-\nu-\mu$. It remains to define $\underline{i}_{\alpha},\bar{i}_{\alpha},\underline{i}_{\beta},\bar{i}_{\beta},\underline{j}_{\alpha},\bar{j}_{\alpha},\underline{j}_{\beta},\bar{j}_{\beta}$. These bounds can be computed from the conditions they satisfy as follows.
$0\leq i_{\alpha}\leq\nu,\ 0\leq i_{\beta}\leq\mu,\ i_{\alpha}+i_{\beta}\leq i,$ (17)
$i_{\gamma}=i-i_{\alpha}-i_{\beta}\leq N-\nu-\mu=\lambda\ \Leftrightarrow\ i-\lambda\leq i_{\alpha}+i_{\beta}$ (18)
From (17) we have $i_{\alpha}\leq\nu$ and $i_{\alpha}\leq i-i_{\beta}\leq i$. From (18) we have
$i_{\alpha}\geq i-\lambda-i_{\beta}\geq i-\lambda-\mu=i+\nu-N.$
We obtained $\max(0,i+\nu-N)\leq i_{\alpha}\leq\min(i,\nu)$ and hence,
$\bar{i}_{\alpha}=\min(i,\nu),\ \underline{i}_{\alpha}=\max(0,i+\nu-N).$
Next, from (17) we have $i_{\beta}\leq\min(i-i_{\alpha},\mu)$. Frow (18) it follows $i_{\beta}\geq\max(0,i-i_{\alpha}-\lambda)$. Hence, $\bar{i}_{\beta}=\min(i-i_{\alpha},\mu)$ and $\underline{i}_{\beta}=\max(0,i-i_{\alpha}-\lambda)$. Further, we have
$0\leq i_{\alpha}+j_{\alpha}\leq\nu,\ 0\leq i_{\beta}+j_{\beta}\leq\mu,\ j_{\alpha}+j_{\beta}\leq j.$ (19)
$j_{\gamma}=j-j_{\alpha}-j_{\beta}\leq N-\nu-\mu-(i-i_{\alpha}-i_{\beta})=\lambda-i_{\gamma}.$ (20)
From (19) it follows $j_{\alpha}\leq\min(j,\nu-i_{\alpha})$. From (20) and (19) we obtain
$j_{\alpha}\geq j-j_{\beta}-(\lambda-i_{\gamma})\geq j-(\mu-i_{\beta})-(\lambda-i_{\gamma}).$
Hence, $\bar{j}_{\alpha}=\min(j,\nu-i_{\alpha})$, $\underline{j}_{\alpha}=j-(\mu-i_{\beta})-(\lambda-i_{\gamma}).$ Now it remains to find $\bar{j_{\beta}}$ and $\underline{j}_{\beta}$. From (19) we have
$j_{\beta}\leq\min(j-j_{\alpha},\mu-i_{\beta}).$
From (20) it follows $j_{\beta}\geq\max(0,j-j_{\alpha}-(\lambda-i_{\gamma}))$. Hence,
$\bar{j}_{\beta}=\min(j-j_{\alpha},\mu-i_{\beta}),\ \underline{j}_{\beta}=\max(0,j-j_{\alpha}-(\lambda-i_{\gamma})).$
Finally, we obtain
${\bf q}_{\nu\mu}$ $=$ $\sum_{i=0}^{n}\sum_{j=0}^{m}\frac{{\bf c}_{ij}}{\binom{N}{i,j}}\sum_{i_{\alpha}=\underline{i}_{\alpha}}^{\bar{i}_{\alpha}}\sum_{i_{\beta}=\underline{i}_{\beta}}^{\bar{i}_{\beta}}\sum_{j_{\alpha}=\underline{j}_{\alpha}}^{\bar{j}_{\alpha}}\sum_{j_{\beta}=\underline{j}_{\beta}}^{\bar{j}_{\beta}}$
$\binom{\nu}{i_{\alpha}}\binom{\mu}{i_{\beta}}\binom{\lambda}{i_{\gamma}}\binom{\nu-i_{\alpha}}{j_{\alpha}}\binom{\mu-i_{\beta}}{j_{\beta}}\binom{\lambda-i_{\gamma}}{j_{\gamma}}a_{1}^{i_{\alpha}}b_{1}^{i_{\beta}}c_{1}^{i_{\gamma}}a_{2}^{j_{\alpha}}b_{2}^{j_{\beta}}c_{2}^{j_{\gamma}},$
$\nu=0,\dots,N,\ \mu=0,\dots,N-\nu,$
where
$\begin{array}[]{l}N=n+m,\ \lambda=N-\nu-\mu,\ i_{\gamma}=i-i_{\alpha}-i_{\beta},\ j_{\gamma}=j-j_{\alpha}-j_{\beta},\\ \underline{i}_{\alpha}=\max(0,i+\nu-N),\ \bar{i}_{\alpha}=\min(i,\nu),\\ \underline{i}_{\beta}=\max(0,i-i_{\alpha}-\lambda),\ \bar{i}_{\beta}=\min(i-i_{\alpha},\mu),\\ \underline{j}_{\alpha}=\max(0,j-(\mu-i_{\beta})-(\lambda-i_{\gamma})),\ \bar{j}_{\alpha}=\min(j,\nu-i_{\alpha}),\\ \underline{j}_{\beta}=\max(0,j-j_{\alpha}-(\lambda-i_{\gamma})),\ \bar{j}_{\beta}=\min(j-j_{\alpha},\mu-i_{\beta}).\end{array}$
$\Box$
Remark 1
The cardinality of the set $I(i_{\alpha},i_{\beta},j_{\alpha},j_{\beta})$ in (16) can also be presented in an equivalent way as follows.
$|I(i_{\alpha},i_{\beta},j_{\alpha},j_{\beta})|=\binom{\nu}{j_{\alpha}}\binom{\mu}{j_{\beta}}\binom{N-\nu-\mu}{j_{\gamma}}\binom{\nu-j_{\alpha}}{i_{\alpha}}\binom{\mu-j_{\beta}}{i_{\beta}}\binom{N-\nu-\mu-j_{\gamma}}{j_{\gamma}}$ (21)
3 Examples and Results
We have implemented and tested the obtained formulae using the Mathematica package (Wolfram Research, Inc., ). In this section we present the results from our experimental work.
Example 1
We consider the parametric polynomial surface $S(u,v)=\left(x(u,v),y(u,v),z(u,v)\right)$ where the coordinate functions $x(u,v),y(u,v),z(u,v)$ are as follows:
$\begin{array}[]{lcl}x(u,v)&=&3u+u^{3},\\[4.30554pt] y(u,v)&=&\frac{7}{5}v+\frac{9}{5}uv-\frac{9}{5}u^{2}v+\frac{13}{5}u^{3}v+\frac{3}{5}v^{2}-\frac{24}{5}uv^{2}+\frac{39}{5}u^{2}v^{2}-\frac{23}{5}u^{3}v^{2}\\[4.30554pt] z(u,v)&=&3u-\frac{15}{4}u^{2}+\frac{19}{20}u^{3}-v-3uv+\frac{117}{10}u^{2}v-\frac{38}{5}u^{3}v+\frac{1}{5}v^{2}+\frac{27}{5}uv^{2}\\[4.30554pt] &&-\frac{141}{10}u^{2}v^{2}+\frac{179}{20}u^{3}v^{2}.\end{array}$
The surface $S$ is represented as $S(u,v)=\sum_{i=0}^{n}\sum_{j=0}^{m}{\bf c}_{ij}u^{i}v^{j}$ where $n=3$ and $m=2$. The corresponding coefficients ${\bf c}_{ij}$, $i=0,\dots,3$, $j=0,\dots,2$, are shown in Table 1. We applied formula (9) for $a=0$, $b=1$, $c=0$, $d=1$ and obtained TPB surface $S_{1}$ of degree $(3,2)$. The control points of $S_{1}$ are shown in Table 1. The surface $S_{1}$ and its control polygon are shown in Figure 1(a).
Further, we applied formula (9) for $a=1/3$, $b=2/3$, $c=1/4$, $d=3/4$ and obtained the TPB surface $S_{2}$ of degree $(3,2)$. The control points of $S_{2}$ are shown in Table 1. Both surfaces $S_{1}$ and $S_{2}$ with their control polygons are shown in Figure 1(b).
Figure 1: Subdivision of TPB surfaces Example 1 using formula (9): (a) The TPB surface $S_{1}(u,v)$ defined for $0\leq u\leq 1$, $0\leq v\leq 1$. (b) $S_{1}$ and the TPB surface $S_{2}(u,v)$ defined for $\frac{1}{3}\leq u\leq\frac{2}{3}$, $\frac{1}{4}\leq v\leq\frac{3}{4}$.
Table 1: The coefficients of surface $S$ and the control points of TPB surfaces $S_{1}$ and $S_{2}$ from Example 1
$S(u,v)$ ${\bf c}_{00}=(0,0,0)$ ${\bf c}_{01}=(0,\frac{7}{5},-1)$ ${\bf c}_{02}=(0,\frac{3}{5},\frac{1}{5})$
${\bf c}_{10}=(3,0,3)$ ${\bf c}_{11}=(0,\frac{9}{5},-3)$ ${\bf c}_{12}=(0,-\frac{24}{5},\frac{27}{5})$
${\bf c}_{20}=(0,0,-\frac{15}{4})$ ${\bf c}_{21}=(0,-\frac{9}{5},\frac{117}{10})$ ${\bf c}_{22}=(0,\frac{39}{5},-\frac{141}{10})$
${\bf c}_{30}=(1,0,\frac{19}{20})$ ${\bf c}_{31}=(0,\frac{13}{5},-\frac{38}{5})$ ${\bf c}_{32}=(0,-\frac{23}{5},\frac{179}{20})$
$S_{1}(u,v),$ ${\bf p}_{00}=(0,0,0)$ ${\bf p}_{01}=(0,\frac{7}{10},-\frac{1}{2})$ ${\bf p}_{02}=(0,2,-\frac{4}{5})$
${\bf p}_{10}=(1,0,1)$ ${\bf p}_{11}=(1,1,0)$ ${\bf p}_{12}=(1,1,1)$
$0\leq u\leq 1$ ${\bf p}_{20}=(2,0,\frac{3}{4})$ ${\bf p}_{21}=(2,1,\frac{6}{5})$ ${\bf p}_{22}=(2,2,\frac{3}{4})$
$0\leq v\leq 1$ ${\bf p}_{30}=(4,0,\frac{1}{5})$ ${\bf p}_{31}=(4,2,\frac{1}{4})$ ${\bf p}_{32}=(4,3,\frac{3}{4})$
$S_{2}(u,v),$ ${\bf p}_{00}=(\frac{28}{27},\frac{983}{2160},\frac{3637}{8640})$ ${\bf p}_{01}=(\frac{28}{27},\frac{385}{432},\frac{781}{2880})$ ${\bf p}_{02}=(\frac{28}{27},\frac{901}{720},\frac{2701}{8640})$
${\bf p}_{10}=(\frac{38}{27},\frac{527}{1080},\frac{613}{1080})$ ${\bf p}_{11}=(\frac{38}{27},\frac{205}{216},\frac{7}{15})$ ${\bf p}_{12}=(\frac{38}{27},\frac{469}{360},\frac{571}{1080})$
$\frac{1}{3}\leq u\leq\frac{2}{3}$ ${\bf p}_{20}=(\frac{49}{27},\frac{1157}{2160},\frac{275}{432})$ ${\bf p}_{21}=(\frac{49}{27},\frac{451}{432},\frac{431}{720})$ ${\bf p}_{22}=(\frac{49}{27},\frac{1039}{720},\frac{1399}{2160})$
$\frac{1}{4}\leq v\leq\frac{3}{4}$ ${\bf p}_{30}=(\frac{62}{27},\frac{1321}{2160},\frac{265}{432})$ ${\bf p}_{31}=(\frac{62}{27},\frac{515}{432},\frac{449}{720})$ ${\bf p}_{32}=(\frac{62}{27},\frac{1187}{720},\frac{1469}{2160})$
Example 2
We consider the parametric polynomial surface $S$ defined in Example 1. First, we applied formula (14) for the triangle $\triangle\bf{abc}$ with vertices ${\bf a}=(0,0)$, ${\bf b}=(1,0)$, ${\bf c}=(0,1)$ and obtained the TB surface $S_{3}$ of total degree $5$. The 21 control points of $S_{3}$ are shown in Table 2. The surface $S_{3}$ and its control polygon are shown in Figure 2(a). The TPB surface $S_{1}$ and the TB surface $S_{3}$ with its control polygon are shown in Figure 2(b).
Further, we applied formula (14) for the triangle with vertices ${\bf a}=(0,\frac{1}{2})$, ${\bf b}=(\frac{1}{2},0)$, ${\bf c}=(\frac{1}{2},\frac{1}{2})$ and obtained the TPB surface $S_{4}$ of degree $5$. The 21 control points of $S_{4}$ are shown in Table 2. The TB surface $S_{4}$ and its control polygon are shown in Figure 2(c). Both surfaces $S_{3}$ and $S_{4}$ with its control polygon are shown in Figure 2(d).
Figure 2: Subdivision of TB surfaces from Example 2 using formula (14): (a) The TB surface $S_{3}$ of total degree 5 defined in the triangle with vertices ${\bf a}=(0,0)$, ${\bf b}=(1,0)$, ${\bf c}=(0,1)$. (b) The TPB surface $S_{1}$ and the TB surface $S_{3}$ with its control polygon. (c) The TB surface $S_{4}$ of total degree 5 defined in the triangle with vertices ${\bf a}=(0,\frac{1}{2})$, ${\bf b}=(\frac{1}{2},0)$, ${\bf c}=(\frac{1}{2},\frac{1}{2})$ and its control polygon. (d) The TB surfaces $S_{3}$ and $S_{4}$ with its control polygon.
Table 2: The control points of TB surfaces $S_{3}$ and $S_{4}$ of degree 55 from Example 2
$S_{3}$ ${\bf q}_{05}=(0,0,0)$ ${\bf q}_{04}=(\frac{3}{5},0,\frac{3}{5})$ ${\bf q}_{14}=(0,\frac{7}{25},-\frac{1}{5})$
${\bf q}_{03}=(\frac{6}{5},0,\frac{33}{40})$ ${\bf q}_{13}=(\frac{3}{5},\frac{37}{100},\frac{1}{4})$ ${\bf q}_{23}=(0,\frac{31}{50},-\frac{19}{50})$
${\bf a}=(0,0)$ ${\bf q}_{02}=(\frac{19}{10},0,\frac{77}{100})$ ${\bf q}_{12}=(\frac{6}{5},\frac{2}{5},\frac{143}{200})$ ${\bf q}_{22}=(\frac{3}{5},\frac{16}{25},\frac{1}{10})$
${\bf b}=(1,0)$ ${\bf q}_{32}=(0,\frac{51}{50},-\frac{27}{50})$ ${\bf q}_{01}=(\frac{14}{5},0,\frac{53}{100})$ ${\bf q}_{11}=(\frac{19}{10},\frac{1}{2},\frac{91}{100})$
${\bf c}=(0,1)$ ${\bf q}_{21}=(\frac{6}{5},\frac{4}{5},\frac{103}{200})$ ${\bf q}_{31}=(\frac{3}{5},\frac{81}{100},\frac{3}{20})$ ${\bf q}_{41}=(0,\frac{37}{25},-\frac{17}{25})$
${\bf q}_{00}=(4,0,\frac{1}{5})$ ${\bf q}_{10}=(\frac{14}{5},\frac{4}{5},\frac{11}{20})$ ${\bf q}_{20}=(\frac{19}{10},\frac{9}{10},\frac{219}{200})$
${\bf q}_{30}=(\frac{6}{5},\frac{6}{5},\frac{9}{40})$ ${\bf q}_{40}=(\frac{3}{5},\frac{22}{25},\frac{2}{5})$ ${\bf q}_{50}=(0,2,-\frac{4}{5})$
$S_{4}$ ${\bf q}_{05}=(0,\frac{17}{20},-\frac{9}{20})$ ${\bf q}_{04}=(\frac{3}{10},\frac{31}{50},-\frac{17}{200})$ ${\bf q}_{14}=(\frac{3}{10},\frac{41}{50},-\frac{33}{200})$
${\bf q}_{03}=(\frac{3}{5},\frac{81}{160},\frac{303}{1600})$ ${\bf q}_{13}=(\frac{3}{5},\frac{523}{800},\frac{43}{320})$ ${\bf q}_{23}=(\frac{3}{5},\frac{653}{800},\frac{27}{320})$
${\bf a}=(0,\frac{1}{2})$ ${\bf q}_{02}=(\frac{73}{80},\frac{601}{1600},\frac{2963}{6400})$ ${\bf q}_{12}=(\frac{73}{80},\frac{857}{1600},\frac{2419}{6400})$ ${\bf q}_{22}=(\frac{73}{80},\frac{221}{320},\frac{2051}{6400})$
${\bf b}=(\frac{1}{2},0)$ ${\bf q}_{32}=(\frac{73}{80},\frac{269}{320},\frac{1859}{6400})$ ${\bf q}_{01}=(\frac{5}{4},\frac{87}{400},\frac{253}{400})$ ${\bf q}_{11}=(\frac{5}{4},\frac{2}{5},\frac{3551}{6400})$
${\bf c}=(\frac{1}{2},\frac{1}{2})$ ${\bf q}_{21}=(\frac{5}{4},\frac{459}{800},\frac{1593}{3200})$ ${\bf q}_{31}=(\frac{5}{4},\frac{591}{800},\frac{2953}{6400})$ ${\bf q}_{41}=(\frac{5}{4},\frac{179}{200},\frac{713}{1600})$
${\bf q}_{00}=(\frac{13}{8},0,\frac{109}{160})$ ${\bf q}_{10}=(\frac{13}{8},\frac{87}{400},\frac{503}{800})$ ${\bf q}_{20}=(\frac{13}{8},\frac{679}{1600},\frac{3767}{6400})$
${\bf q}_{30}=(\frac{13}{8},\frac{993}{1600},\frac{3589}{6400})$ ${\bf q}_{40}=(\frac{13}{8},\frac{129}{160},\frac{349}{640})$ ${\bf q}_{50}=(\frac{13}{8},\frac{157}{160},\frac{347}{640})$
4 Conclusions
Blossoming is an important tool in the analysis and manipulation of polynomial and piecewise polynomial curves and surfaces, significantly enhancing the capabilities within theoretical and applied CAGD . This paper addresses the challenge of deriving control points for subdivided Bézier curves and surface patches via blossoming. The proposed closed-form formulae computationally optimize the evaluation process for subdivision of Bézier curves, and TPB and TB surfaces. We believe that the presented evaluation process based on closed-form formulae can be significantly optimized further by providing a more efficient computational procedure computing control points in bulk instead of computing each control point separately. Another direction of research is to consider blossoming approach in non-polynomial cases and look for speed-up computational procedures for evaluation of the corresponding blossoms.
Acknowledgments
This research was supported in part by Sofia University Science Fund Grant No. 80-10-11/2025 and the project UNITe BG16RFPR002-1.014-0004 funded by PRIDST.